### v2

@@ Abstract
In this study, we address the challenge of detecting “discursive communities” on X/Twitter by focusing on the role of verified users as the main content creators in online political debates. The analysis centers on three major Italian political events in 2022 – the Presidential election, a governmental crisis, and the general elections – occurring before the introduction of paid account verification. We propose and compare two novel methodologies, MonoDC and BiDC, which exploit, respectively, the retweet network among users and a similarity network based on shared audiences, while integrating a maximum entropy null model to filter out the inherent noise in online social networks. Our results demonstrate that leveraging verified users—considered as indicators of prestige and authority—leads to significantly clear community partitions that closely reflect the actual political affiliations, outperforming standard community detection algorithms applied to the entire retweet network. Moreover, the comparison of different methodologies and user sets suggests that the status conferred by the blue verification tick plays a dominant role in shaping online discourse, with important implications for platform governance, especially in light of the recent shift to paid verification.
1 Introduction
Online communities naturally emerge from users’ tendency to follow and interact with others who share similar interests, experiences, connections, or political views [1, 2, 3]. Such communities can provide valuable support and shared knowledge [4, 5, 6]; however, they can also form “echo chambers,” where members are exposed only to a narrow range of perspectives that reinforce their existing beliefs [7, 8, 9, 10, 11, 12, 13]. To fully understand the influence of user communities on information diffusion and opinion dynamics, it is crucial to develop effective methods for detecting these communities and capturing their key structural features.
In the literature, several refined approaches have been proposed to enhance standard community detection methods by tailoring them to the context of online social networks. For example, Citraro and Rossetti [14] incorporate account attributes to strengthen the conventional definition of communities. Other approaches leverage the varied roles played by different accounts in online debates. Although Twitter/X was initially designed as a many-to-many platform, numerous studies have observed that it actually functions as a few-to-many network, where a small number of influential content creators generate the majority of posts [15, 16, 17]. In this context, it becomes natural to focus on the leading content creators for detecting communities on Twitter/X.
Identifying which accounts should be considered as leading content creators is, therefore, essential. In [4], influential users—defined as those with more followers than followees and with more mentions received than given—are identified as the primary figures in the debate. Similar approaches, based on user activity, have also been explored [15, 16, 17]. Other studies place more emphasis on account prestige than on activity, considering verified users as the leading content creators [18, 19, 20].
Verified users on X/Twitter have an interesting history: they were first introduced in June 2009 to curb the impersonation of notable persons or organizations [21, 22]. In this early stage, the platform assumed responsibility for identifying key public figures and verifying their identities with a blue checkmark, while standard users were not part of the verification process [23]. Efforts to broaden the pool of verified accounts through user applications were unsuccessful [21]. Eventually, in November 2022, verification became a feature of X’s Premium accounts, available upon payment [24]—a change that has sparked concerns regarding the influence of verified users on online debates [25, 26].
In the present study, we investigate the method proposed in [18] using three distinct datasets collected during 2022, in a period before blue ticks could be purchased via X’s Premium subscription. Our framework comprises: (i) identifying verified users as the leading content creators; (ii) constructing an informative network of verified users by filtering out random noise through a comparison between empirical data and an appropriate null model; (iii) detecting communities within this sub-network; and (iv) propagating the labels to all other users. Following the terminology introduced by [27], the resulting partition is referred to as “discursive” communities.
A key component of our approach, in addition to leveraging verified users, is the elimination of random noise from the system by comparing empirical data against a maximum entropy statistical benchmark. As numerous studies have shown, removing random noise is crucial in online social networks to extract the relevant signal regarding users’ preferences and behaviors [18, 28, 29]. In this regard, maximum entropy null models provide an ideal statistical benchmark, being maximally random apart from the constraints imposed [30].
@@ 1 Introduction
Figure 1: MonoDC’s pipeline. Starting from the data from Twitter/X, we build a bipartite directed network between accounts and posts; a directed arrow from a user to a post represents authorship, while in the opposite direction represents a retweet. Then, the content of the bipartite network is projected into a monopartite directed network of users. Such a network is further statistically validated using the expectations of a maximum entropy null model for bipartite directed networks. Then, the sub-network relative to content creators is analysed, and communities are detected using standard algorithms. Finally, the so-obtained labels are propagated to standard users through a label propagation algorithm.
Compared to previous work, we distinguish between two main versions of the algorithm that differ in how step (ii) is implemented. In $\mathrm{MonoDC}$ (Monopartite-network-induced Discursive Communities), we consider the direct interactions —i.e., retweets— among content creators; a cartoon representation of the various steps of the algorithm is displayed in Fig. 1. In $\mathrm{BiDC}$ (Bipartite-network-induced Discursive Communities), we instead construct a similarity network based on the shared audience between pairs of leading content creators; a cartoon representation of the various steps of the algorithm is displayed in Fig. 2. Additionally, we explore various classes of leading content creators beyond verified users.
Our proposed methods were tested on three large datasets linked to the primary political events in Italy during 2022, before user verification became a feature of Premium accounts. To validate our results, we manually annotated approximately 300 verified accounts to determine their political leanings and compared different clustering algorithms to assess their ability to detect this signal.
The key finding is that both $\mathrm{MonoDC}$ and $\mathrm{BiDC}$, when using verified users as the main content creators, significantly outperform standard algorithms. They produce clearer partitions that consist of a few meaningful communities, which accurately reflect the political landscape during the studied period. Specifically, $\mathrm{MonoDC}$ appears to capture political parties more effectively—likely because politicians predominantly retweet their party leaders—while $\mathrm{BiDC}$ better identifies political alliances due to the shared construction of narratives. Importantly, our results deteriorate markedly when the set of main content creators is altered.
@@
Figure 2: BiDC’s pipeline. Starting from the data from Twitter/X, we build a bipartite undirected network between the accounts of the leading content creators and the one of standard users; a link is present if the standard users retweeted the given content creator at least once. Then, the information of the bipartite network is projected into a monopartite network of leading content creators. Such a network is further statistically validated using the expectations of a maximum entropy null model for bipartite undirected networks. Finally, the so-obtained labels are propagated to standard users through a label propagation algorithm.
In summary, our study demonstrates that verified accounts play a central role in shaping discursive communities on X/Twitter. This evidence carries important implications, especially considering that our datasets are limited to the period before the implementation of X’s Premium accounts, when blue checkmarks were granted based on the platform’s autonomous verification of key accounts. While we cannot conclusively determine whether the effectiveness of verified users as leading content creators stems solely from the prestige of the verification or from other factors, our findings indicate that definitions based exclusively on user activity yield inferior performance. If the prominence of verified users in steering the online discourse is indeed due to the prestige associated with blue ticks, then introducing verification as a paid feature could pose serious challenges to the platform’s integrity, potentially enabling malevolent actors to purchase authority to distort the political agenda.
The paper is organised as follows. In Section 2, we introduce the methodology adopted in this study. Specifically, we present maximum entropy null models and their application to the detection of discursive communities (Subsections 2.1 and 2.2). We also define the V-measure [31], the metric used to compare different partitions (Subsection 2.3). Section 3 presents our results. We begin by illustrating the limitations of standard methods in reproducing non-trivial benchmarks observed in real-world data (Subsection 3.3). We then show how $\mathrm{MonoDC}$ and $\mathrm{BiDC}$ address these issues when verified users are identified as the primary content creators (Subsection 3.4). Additionally, we show that the performance of both $\mathrm{MonoDC}$ and $\mathrm{BiDC}$ deteriorates when different groups of leading content creators are considered (Subsection 3.5). Finally, in Section 4, we summarise our findings and discuss their potential implications.
2 Methods
2.1 Maximum-entropy null models and discursive communities
As underlined by many studies across a multitude of applications  [18, 19, 20, 27, 32, 33, 34, 35, 36, 28, 29, 13], using a proper statistical validation is critical for the analysis of online social networks (OSNs) that are particularly prone to noise. Entropy-based null models [30] represent a tailored and, at the same time, unbiased tool for detecting the statistically significant signal in complex networks. The method essentially consists in: i) defining an ensemble of graphs, i.e. the set of all possible graphs having the same number of nodes as the observed network; ii) performing a constrained maximization of the Shannon entropy of the ensemble, fixing some constitutive properties of the original system [37]; iii) maximizing the (log-)likelihood of the observed network [38], to get the numerical values of the parameters of an effective fitness model.Remarkably, this framework mimics Jaynes’ derivation of statistical ensembles in Physics starting from Information Theory [39].
Let us start from a real network $G^{*}$ (in the following, all quantities related to the observed network will be denoted by an asterisk *). Firstly, let us define a set $\mathcal{G}$ – the ensemble– including all possible graphs having the same number of nodes $N$ as in $G^{*}$ [^1] Since the number of nodes is fixed in the randomization, we drop the asterisk * from $N$.. Let us identify a vector of quantities $\vec{C}$ that are fundamental for the description of the real system. A proper benchmark for $G^{*}$ would be completely random, but for the information contained in $\vec{C}(G^{*})$: in this way it is possible to highlight all properties of $G^{*}$ that are not trivially due to $\vec{C}$. Translating this rationale in maths, we need to perform a constrained maximization of the Shannon entropy (to have a maximally random benchmark) associated to the ensemble, where the constraints are exactly $\vec{C}(G^{*})$. In this sense, the (ensemble) average $\langle\vec{C}\rangle$ will be set exactly to $\vec{C}(G^{*})$.
If $S$ is the Shannon entropy, defined as
$S=-\sum_{G\in\mathcal{G}}P(G)\ln P(G),$
$P(G)=\frac{e^{-\vec{\theta}\cdot\vec{C}(G)}}{Z(\vec{\theta})},$ (1)
$\langle\vec{C}\rangle=\vec{C}(G^{*}).$ (2)
It can be shown that the condition in Eq. (2) can also be obtained through the maximization of the (log-)likelihood [38].
The framework described above is particularly flexible and can be adapted to different kinds of networks. In the present paper, we are interested in its bipartite versions for bipartite undirected and bipartite directed networks, i.e. respectively, the Bipartite Configuration Model (or BiCM [40]) and the Bipartite Directed Configuration Model (or BiDCM [41, 18]). In both cases, the constraints imposed in the entropy maximisation are, respectively, the undirected and directed degree sequences. More details about BiCM and BiDCM can be found in Appendices A.1 and A.2.
2.2 $\mathrm{MonoDC}$ and $\mathrm{BiDC}$
In the following, the two methods that we are proposing are presented. They both leverage the presence of leading content creators, i.e. users mostly posting original messages and less inclined to repost the messages created by others. There are multiple ways to define leading content creators: at the present level, we prefer to remain as general as possible. Therefore, in the following those users are called $\top$-⊤ -users, while the $\top$ users are called $\bot$-⊥ -users.
2.2.1 $\mathrm{MonoDC}_{\top}$
At least in the Italian political debate, it is common to observe that less relevant political figures are particularly active in liking and reposting the content created by more central leaders [^2] https://www.ilpost.it/2024/10/31/post-giorgia-meloni-social-commenti-parlamentari-fdi/. There could be multiple reasons for such a behaviour: intercept the audience of the leaders, demonstrate loyalty to either the leaders, the party or the objective of the campaign, or simply to contribute to the campaign of the leaders by increasing the visibility of their content. Therefore, it is reasonable to expect the activity of politicians to be much more focused on the activity of their party than other users.
To capture such a phenomenon, the MonoDC studies the interactions among content creators through their activity in creating new contents or reposting the one created by others. The steps of MonoDC are represented as a cartoon in Fig. 1. We describe data from Twitter/X as a bipartite directed network between users and posts in which edges from users to posts indicate authorships, while the vice versa indicates retweets. The information contained in the directed bipartite network is then projected to describe the retweet network, i.e. a directed multilink network in which nodes are users and the number of edges from user $A$ to user $B$ is the number of retweets that $B$ made to the post written by $A$. In this representation, we expect the (bipartite directed) degrees to play a role, since they represent the activity of users in either creating new posts or retweeting existing ones (represented, respectively, as the out- and in-degree in the bipartite directed network of the user nodes) and the virality of the various posts (i.e. the out-degree of the post nodes). Therefore, we can use a maximum entropy null model for directed bipartite networks as a statistical benchmark to validate the empirical retweet network, i.e. the Bipartite Directed Configuration Model (BiDCM) mentioned in Section 2.1. Technical details about BiDCM can be found in Appendix A.2.
The result of such validation is a directed binary network, in which an edge represents that the retweet activity between the two nodes is non-trivial, i.e. it cannot be explained by just the attitude of users towards posting and reposting and the virality of the messages shared. We further restrict the validated network to only users in the set of content creators, i.e. the $\top$ user, in our jargon.
The main idea is that the null model is not informed about the orientation of the users and therefore expects the debate to be “flat”, i.e. without any particular focus of the users towards the content created by some of the accounts. In this sense, communities should emerge naturally by comparing empirical data with the null model. Such a structure can be further captured by running a community detection algorithm on the validated retweet network.
Finally, the labels obtained by the community detection are extended to $\bot$ users (i.e. the users that are not in the $\top$ set) by implementing Raghavan’s Label propagation algorithm [42] on the retweet network, using as label seeds the ones obtained by the community detection run on the validated retweet network restricted to $\top$-users.
2.2.2 $\mathrm{BiDC}_{\top}$
The intuition behind $\mathrm{BiDC}_{\top}$ is to capture discursive communities through the perception of “standard” users of the similarity in the content created by the leaders of the debates [18, 43, 27, 19]. Such a phenomenon is more stressed during political debates, being more or less intense depending on the electoral system of the country [43].
Following the procedure outlined in Fig. 2, we built the bipartite network between leading content creators and “standard” users: a link is present if the standard user has retweeted at least once the corresponding content creator. We then projected the bipartite network to a monopartite co-occurrences network among content creators. As in the case of $\mathrm{MonoDC}_{\top}$, we validate the network by comparing the empirical data with a maximum entropy null model for bipartite networks, i.e the Bipartite Configuration Model (BiCM) mentioned in Section 2.1. In this case, the information about the degree of the users is capturing the number of different retweeters and the number of different content creators retweeted for $\top$-⊤ - and $\bot$-⊥ -users, respectively; more details about the implemented maximum entropy null model can be found in Appendix A.1.
BiCM, while accounting for the degree sequences of both layers, is not informed about the presence of communities and therefore expects the co-occurrence network not to have a modular structure. In this sense, any non-trivial focus of the users towards groups of content creators should be captured by comparing empirical data against the null model. Such a structure can be further refined by running a community detection algorithm on the validated projection network.
Similarly to $\mathrm{MonoDC}_{\top}$, labels are extended to $\bot$ users by using Raghavan’s Label propagation algorithm [42] on the retweet network, using as label seeds the ones obtained by the community detection run on the validated retweet network.
2.3 Measuring the similarity of two partitions
To assess the similarity between two ways of partitioning the same set of users, we use of the V-measure ($\mathrm{VM}_{\beta}$), a parametric measure of cluster similarity proposed by [31].
$\mathrm{VM}_{\beta}(P,Q)=\frac{(1+\beta)I(P,Q)}{S(P)+\beta S(Q)}$
where $I(P,Q)$ is the mutual information between $P$ and $Q$.
For $\beta=1$ we have
i.e., the $\mathrm{VM}_{1}$ accounts for the mutual information between the two partitions in a way that does not depend on their entropy, being equivalent to the Variation of Information [44], normalized by the total entropy of the partitions. With respect to other information-theoretic metrics, the $\mathrm{VM}_{1}$ and the VI have the advantage of being two distances – in particular, of satisfying the triangle inequality. Further, the $\mathrm{VM}_{1}$ always takes values in $[0,1]$, making it possible to directly compare the scores obtained for different combinations of algorithms and/or datasets. By letting $\beta$ vary in $[0,+\infty)$, we consider both scenarios in which one partition consists of sub-communities (e.g., political movements/trends within parties) or super-communities (e.g., political coalitions) of the other one. More details about the V-measure can be found in Appendix B.
3 Results
3.1 Data and networks
To compare our approach for discursive community with standard community detection algorithms, we collected Twitter data regarding the three main political events debated on Italian Twitter in 2022. For each such topic of discussion, we identified a set of relevant keywords that were used to collect tweets using Twitter’s Academic API. Let us remark that data were downloaded before user’s verification became a feature of the X’s Premium subscription [^3] https://en.wikipedia.org/wiki/Twitter_verification.. A summary of the datasets considered in this paper is reported in Table 1.
@@ 3.1 Data and networks
Dataset Unique users Unique tweets Annotated users Verified users Influential users H-$\geq$3 users
@@
3.2 The benchmark
In general, there is no reason to expect data to cluster due to some metadata [45]. Multiple effects can cause –especially in online social networks– different users to interact with the same content creators. Nevertheless, due to the strictly political nature of the debates captured by our data, we expect to observe a signal of the political orientation of the users in the cluster that can be used as a benchmark for the classification methods under study.
To construct our benchmark labeling, we first focused on just the verified accounts of Italian politicians and political parties. We choose verified accounts since their identities are certified by the platform. We manually labeled these accounts according to their political party and then grouped these labels according to the coalition each party belonged to during the general elections. This procedure yielded two annotation levels, which we refer to as “parties” and “coalitions”.
To check if this benchmark is well-posed, we studied the retweet network restricted to the annotated users. Retweets are usually considered as a sort of endorsement of the content shared [1], so we expected users with a similar political orientation to cluster in this network. The modularity of the partitions induced by the annotations, whose values are all above 0.6 (cfr. Table 2), confirms our intuition and supports the idea of measuring the quality of a community detection algorithm also based on its ability to retrieve our manual annotations.

### v3

@@ Abstract
Online discussions are often characterized by strong behavioral asymmetries: a relatively small fraction of users actively produces content, while the majority primarily consumes and redistributes it. Here we propose a community-detection framework for online social networks that exploits this asymmetry by first identifying and clustering a set of leading users, and then extending the resulting labels to the broader user base. We introduce two complementary strategies to cluster leaders, one based on their mutual interactions and the other on audience overlap, both relying on entropy-based filtering to separate signal from noise. We evaluate the framework on three major Italian political debates on Twitter/X, using public figures–identified through the pre-2022 verification system–as leaders, and known affiliations of political actors as ground truth labels. Compared with standard baselines, the proposed approach yields more coherent and interpretable communities aligned with political structures, with the two variants respectively recovering parties and coalitions. Activity-based criteria for selecting leaders produce qualitatively similar but consistently weaker results, particularly at the coalition level. Overall, our findings show that creating statistically validated networks of publicly recognized figures, whose off-platform roles constrain and stabilize their online behavior, provide a strong basis to identify discursive communities on social media. Although developed for Twitter/X, the approach is conceptually general, as it leverages structural asymmetries common to many online platforms.
Introduction
Online social platforms have become crucial arenas for public discourse, with communities forming organically around shared narratives, values, and ideological affinities [Conover2011a, Conover2011b, Conover2012]. While these discursive communities can offer solidarity and foster public engagement [GonzalezBailon2013, Andalibi2018, Morini2025], they also risk becoming echo chambers—insulated environments that reinforce existing beliefs and suppress dissenting perspectives [jamieson08echo, Garrett2009, delvicario2016, Zollo2017Debunking, Villa2021, Morini2021, Pratelli2024]. Understanding the structural dynamics that govern these communities is therefore essential to evaluating how information circulates and how opinions crystallize in digital spaces.
Traditional approaches to community detection in social networks have evolved significantly, incorporating network topology, account metadata, and behavioral signals. For instance, some methods enhance classical algorithms by integrating account characteristics [citraro2019eva], while others acknowledge that platforms like Twitter/X function less as egalitarian many-to-many networks, and more as hierarchies dominated by a small number of highly visible content creators [Watts2007, Wu2011, Hilbert2017]. Within this structure, influence is disproportionately concentrated, making the identification of these “elite” accounts central to the analysis of online discourse [Becatti2019d, Caldarelli2020b, Radicioni2021a].
Building on this body of work, the first contribution of this article is the definition of a general framework for detecting online discursive communities, formalizing the idea that online debates are driven by a relatively small set of users who actively produce content. Because content production entails greater exposure and commitment—especially in political discussions—these users tend to adopt more consistent and publicly interpretable positions. Once content producers are identified, the affiliation of less active users can be inferred from their interaction patterns, even when individual behavior is sparse.
The proposed framework operationalizes this rationale in a relatively simple way. We partition the user set $U$ into two disjoint subsets, $U=\top\cup\bot$, cluster the content creators (the set $\top$) by applying a standard community-detection algorithm to an appropriately defined network $G=(\top,E)$, and then extend the resulting labels to users in $\bot$ based on their interactions with $\top$.
A key modeling choice, however, concerns the construction of the edge set $E$, which defines how relationships among content creators are represented. We consider two alternatives: one based on statistically significant interactions directly observed among users in $\top$, and one based on a validated projection of the bipartite network linking content creators in $\top$ to their audiences in $\bot$. These two variants define Monopartite-network-induced Discursive Communities ($\mathrm{MonoDC}_{\top}$) and Bipartite-network-induced Discursive Communities ($\mathrm{BiDC}_{\top}$), respectively. Both methods rely on maximum-entropy null models [Cimini2019] to statistically validate inferred connections and to filter out noise–a well-known challenge in the analysis of online social networks [Becatti2019d, Declerck2022a]. Statistical filtering plays a limited role in $\mathrm{MonoDC}_{\top}$, but is essential in $\mathrm{BiDC}_{\top}$ to remove spurious correlations induced by heterogeneous audience activity.
Beyond the construction of $E$, the framework admits different criteria to identify the set $\top$ of content creators. From this perspective, the second contribution of this article is less methodological and more substantive: we show that, although social-media relevance can be acquired in an anonymous and platform-driven manner, ideological communities in political debates remain predominantly anchored to publicly recognizable political figures.
Accordingly, we focus on Twitter’s verified accounts ($\top=V$), which, at the time covered by our datasets, identified public figures whose activity on the platform was closely tied to their off-platform roles. All datasets analyzed precede the introduction of paid verification in November 2022, and therefore correspond to this original, institutionally grounded meaning of verification [Haman2023].
We compare this choice with two activity-based criteria derived from the data. We consider influential users ($\top=I$) as defined by [GonzalezBailon2013], based on follower structure and mention activity, and high-$h$-index users ($\top=H$), obtained by adapting the Hirsch index [Hirsch2005] to retweet dynamics. These data-driven selections serve as baselines to evaluate the added value of anchoring the analysis on publicly recognized figures.
We analyze three datasets covering major political events in Italy in 2022, where discursive communities are expected to mirror political alignments. When evaluated against the known affiliations of clearly identifiable political actors, $\mathrm{MonoDC}_{V}$ captures party-level structures, whereas $\mathrm{BiDC}_{V}$ is more effective at identifying coalitional configurations. Standard community-detection algorithms applied to the complete retweet network yield noisier and less interpretable partitions, while activity-based criteria perform better than these baselines but remain less reliable than anchoring the analysis on verified accounts, especially at the coalition level.
We argue that this is not an effect of verification, but, again, a consequence of the fact that, historically, only publicly recognizable figures were “verifiable”, a requisite that in turn constrained and stabilized their online behavior. In this sense, our results show that, even in environments where visibility can be gained in largely anonymous and platform-driven ways, ideological communities in political debates remain primarily structured around actors endowed with public recognizability beyond the platform. By foregrounding this mechanism, we propose and analyze a framework that, while developed for Twitter/X, is conceptually general, as it leverages asymmetries between content producers and audiences that characterize many contemporary online platforms.
Finally, we show that a carefully selected subset of users and interactions captures a large fraction of the information needed to reconstruct political communities, offering practical insights for designing efficient data-collection strategies under increasingly restrictive platform access conditions.
Related work
A substantial body of research has investigated the structure of the online discourse on Twitter and other social media platforms by analyzing retweet networks, information cascades, and statistically validated interaction patterns (e.g., [garimella2018quantifying, Becatti2019d, cinelli2021echo, lorenzspreen2023digitalmedia]). These studies have consistently shown, in particular, that political debates on social media exhibit strong polarization and modular structure, often aligned with ideological or partisan divides.
A central methodological challenge in this literature concerns the high level of noise and heterogeneity in user activity. To address this issue, several works have adopted statistical filtering techniques based on null models to extract significant interaction patterns from bipartite or projected networks. In particular, maximum-entropy null models have been widely used to preserve local constraints—such as node degrees—while filtering out correlations that can be explained by random activity alone [Cimini2019, Saracco2017]. These approaches have been shown to uncover meaningful mesoscale structures in retweet, mention, and user-content networks, enabling the identification of latent ideological alignment from interaction data [Becatti2019d, Caldarelli2020b, Radicioni2021a, Mattei2022, Pratelli2024].
Within this line of work, the Italian political context has received special attention. Becatti et al. [Becatti2019d] have shown that entropy-based filtering improves the interpretability and stability of political communities compared to unfiltered network representations (the method proposed therein is called $\mathrm{BiDC}_{V}$ in the present manuscript). Related studies have further demonstrated that statistically validated user–hashtag bipartite projections can reveal well-defined discursive communities during national electoral campaigns, even in the absence of explicit ideological labels [Radicioni2021a].
A parallel strand of research has focused on the role of influential or central users in shaping the online discourse. Early work has defined influence primarily in terms of activity-based or structural measures, such as the number of followers, retweets, mentions, or centrality in interaction networks [Watts2007, Hilbert2017, Wu2011]. Other studies have explored the diffusion power of opinion leaders and elites in online debates, often highlighting the disproportionate impact of a small fraction of highly visible users [bakshy2011everyone].
Despite these advances, two important gaps persist in the literature. First, most community-detection approaches implicitly treat users symmetrically, without explicitly accounting for the strong asymmetry between content producers and content consumers that characterizes participation on social media–in general, and especially for political debates. Second, benchmarks for evaluating communities are often based on internal network metrics, such as modularity or stability, rather than on externally validated ground truth labels. As a result, it remains unclear (i) whether institutional signals of authority–such as pre-2022 verified status on Twitter–provide particularly informative anchors for community detection, and (ii) how alternative criteria for identifying leading users affect the resulting mappings when embedded into statistically validated pipelines. The present work addresses these gaps by proposing a leader-based framework that explicitly separates content creators from their audiences and systematically compares different strategies for selecting politically relevant producers of content.
Methods
We investigate how online discursive communities form around publicly relevant users on Twitter/X by analyzing large-scale interaction data. While the framework we propose is general, we focus empirically on three major political events in Italy during 2022, a domain in which online debates tend to be especially structured and externally grounded. Our approach builds on the observation that social-media debates are typically characterized by strong asymmetries in user behavior: a relatively small subset of users actively produces salient content, whereas the majority primarily consumes and redistributes it. This asymmetry is particularly pronounced in political discussions, where public figures and institutional actors play a central role. To exploit this structure, we adopt a leader-based framework in which community detection is first performed on a selected set of content creators and subsequently extended to the broader user base.
Detection of political discursive communities
To identify meaningful structures in online political debates, it is necessary to distinguish statistically significant interaction patterns from correlations induced by heterogeneous user activity and content popularity. To this end, our approach relies on maximum-entropy null models as a principled statistical baseline for validating inferred network connections before applying community-detection algorithms.
Within this framework, an ensemble of random networks is defined by maximizing the Shannon entropy subject to a set of local constraints that reproduce selected properties of the observed data, such as degree sequences or activity levels [Jaynes1957, park2004statistical, Cimini2019]. Empirical links are evaluated against their expected occurrence under the null model, and only statistically significant deviations are retained. Model parameters are estimated by maximizing the likelihood of the observed network [Garlaschelli2008], and statistical significance is assessed using false discovery rate correction [Benjamini1995].
This maximum-entropy approach is grounded in information theory and statistical mechanics and has proven effective across a wide range of networked systems [Becatti2019d, Caldarelli2020b, Caldarelli2021, Radicioni2021a, Radicioni2021b, Mattei2021, Guarino2021, Mattei2022, Bruno2022, Declerck2022a, Declerck2022b, Pratelli2024]. In the following, we introduce two complementary methods that instantiate this framework in different ways, depending on how interactions among selected content creators are inferred. Further methodological details are reported in Section A of the Appendix.
In the analyses presented here, we focus on verified accounts as a concrete instantiation of content creators, as–prior to the introduction of paid verification–they largely corresponded to publicly recognized figures. However, the methods described below are general and can be applied to any set $\top$ of users identified as debate leaders.
Construction of the $\mathrm{MonoDC}_{V}$ network
The first method, $\mathrm{MonoDC}_{V}$ (Monopartite-network-induced Discursive Communities; Fig. 1), focuses on retweet interactions among verified users. We start by constructing a bipartite directed network between users and tweets, where directed edges represent authorship and retweet actions. At this stage, we consider all verified users together with all users who retweeted at least one verified account in the dataset under study. This structure is then projected onto a weighted directed user-user network, in which a link from user $i$ to user $j$ encodes the number of times $i$ retweeted content authored by $j$.
To separate meaningful interaction patterns from noise–an intrinsic feature of online social activity–we apply a maximum-entropy null model tailored to bipartite directed networks. The model uses the observed posting and retweeting activity of users as constraints while randomizing all other aspects, providing a principled statistical baseline. For each pair of users $i,j$, we compute the probability of the observed retweeting weight under the null model, and retain the edge only if its p-value is statistically significant (more details are present in Section A of the Appendix).
Community detection is then performed on the subnetwork of the validated network induced by verified users. The presented results have been obtained with a modularity-based algorithm, but we verified that alternative algorithms yield essentially the same results. Finally, the detected community labels are propagated to non-verified users using a standard label-propagation scheme [Raghavan2007a].
@@ Methods
Figure 1: $\mathrm{MonoDC}$’s pipeline. Starting from the data from Twitter/X, we build a bipartite directed network between accounts and posts; a directed arrow from a user (a circle) to a post (a square) represents authorship, while in the opposite direction represents a retweet. Then, the content of the bipartite network is projected into a monopartite directed network of users. Such a network is further statistically validated using the expectations of a maximum entropy null model for bipartite directed networks. Then, the sub-network relative to the selected content creators (magenta circles) is analysed, and communities are detected using standard algorithms. Finally, the so-obtained labels are propagated to standard users (blue circles) through a label propagation algorithm. The icon representing Twitter/X data is obtained by superimposing the icons “Data Base #7994662” by Farrih Icon and “Hashtag #4315467” by Alex Burte, both from Noun Projects.
Construction of the BiDCV network
The second method, $\mathrm{BiDC}_{V}$ (Bipartite-network-induced Discursive Communities; Fig. 2[Becatti2019d]), infers relationships among verified users indirectly, based on the similarity of their audiences. Here, we construct a bipartite undirected network connecting verified users to all users who retweeted them at least once during the event.
This bipartite structure preserves the heterogeneity between content creators and their audiences. To extract meaningful similarities between verified users, we apply a maximum-entropy null model for bipartite undirected networks, which accounts for differences in audience sizes and engagement patterns. Statistically significant similarities are retained, yielding a validated projection onto a network of verified users.
Community detection is performed on this projected network, and the resulting labels are propagated to non-verified users according to the composition of their retweet activity. As in $\mathrm{MonoDC}_{V}$, the propagation step allows us to assign community memberships to the entire user set starting from a limited number of leader nodes.
@@
Figure 2: $\mathrm{BiDC}$’s pipeline. Starting from the data from Twitter/X, we build a bipartite undirected network between the accounts of the content creators and the one of standard users; a link is present if the standard users retweeted the given content creator at least once. Then, the information of the bipartite network is projected into a monopartite network of content creators. Such a network is further statistically validated using the expectations of a maximum entropy null model for bipartite undirected networks. Finally, the so-obtained labels are propagated to standard users through a label propagation algorithm. The procedure was first proposed in [Becatti2019d]. The icon representing Twitter/X data is obtained by superimposing the icons “Data Base #7994662” by Farrih Icon and “Hashtag #4315467” by Alex Burte, both from Noun Projects.
Evaluation
We evaluate the quality of the detected communities using the V-measure ($\mathrm{VM}_{\beta}$[rosenberg2007v], an information-theoretic metric that quantifies the agreement between two partitions based on their mutual information. The V-measure is well suited to our setting, as it remains interpretable across different levels of partition granularity and explicitly balances homogeneity–how well each cluster contains only members of a single class–and completeness–whether all members of a given class are assigned to the same cluster–through the parameter $\beta$. For $\beta=1$, $\mathrm{VM}_{1}$ reduces to the Variation of Information [Meila2007], normalized by the total entropy of the partitions. By varying $\beta$ over $0,+\infty)$, we account for both finer-grained and more aggregated political structures, such as political movements/trends within parties or political coalitions. More details can be found in Section B of the Appendix.
As a baseline, we also apply standard community-detection algorithms directly to the complete retweet network. This network is defined as a directed, weighted graph including all active users, where edges represent the number of retweets between pairs of users. Unlike $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$, this representation is not statistically filtered and serves to isolate the contribution of entropy-based validation and leader selection.
Although MonoDCV and BiDCV ultimately rely on Louvain for final partitioning, the comparison with Louvain applied to the complete retweet network serves a different purpose. In our frameworks, Louvain operates after statistically validated filtering, and with an informative initialization provided by verified seeds. Running Louvain directly on the raw network therefore represents a baseline that isolates the contribution of our preprocessing and seed selection. This allows us to quantify the extent to which discursive communities emerge from the raw interaction structure alone versus from statistically validated and leader-anchored representations.
Data
We analyze three datasets covering the Italian presidential election, a government crisis, and the 2022 general elections. Each dataset consists of millions of tweets and hundreds of thousands of users, collected via Twitter’s Academic Application Programming Interface (API) using event-specific keywords; more details about the queries used for the data collection can be found in Section C of the Appendix.. Table 4 summarizes the main characteristics of the datasets.
Among the verified accounts appearing in each dataset, we manually annotated those corresponding to politically active public figures. Annotations were performed at both the party and coalition level and provide an external ground truth against which we assess the performance of our methods.
Manual annotation of verified accounts
For each event, we manually annotated the political affiliation of verified users to construct the ground truth used in our evaluation. Only verified accounts of political parties and politicians were annotated; accounts belonging to journalists, media outlets, institutions, or entertainers were excluded. Each annotated user was assigned both a party label and a coalition label. The party label was determined based on publicly available information from the user’s Twitter profile and Wikipedia page, or from institutional sources such as party websites, parliamentary records and candidate lists. The coalition label was assigned according to the official coalitions recognized during the 2022 Italian electoral cycle.
Annotations were performed by one of the authors, while the remaining authors conducted random spot checks to assess consistency and correctness. Verified accounts with no clear political alignment were included in the construction of both $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$ networks, but were excluded from the ground truth and used only as unlabelled nodes in the evaluation.
@@
[…]
@@
Code and data availability
Preprocessed data and code used for the analysis are available on GitLab at https://gitlab.com/s.guarino/dico_analysis.
Results
Verified accounts provide a strong political signal under the pre-2022 verification regime
To assess the effectiveness of our framework, we focused on three political events that sparked intense debate on Italian Twitter in 2022: the re-election of the President of the Republic, the resignation of Prime Minister Draghi, and the general elections in September. Crucially, these datasets were compiled before the platform introduced paid verification, ensuring that verified accounts retained their original role as signals of institutional recognition rather than self-declared status.
We manually annotated a subset of verified accounts (between 229 and 292 per event) based on their publicly identifiable political affiliations. When restricting the retweet networks to, first, the Greatest Connected Component (GCC), and then to annotated users and evaluating the partition induced by their known party and coalition labels, we observed highly modular structures. Modularity values exceeded 0.6 in all cases (cfr. Table 2), indicating that retweet interactions among verified users are strongly structured along political lines.
These results show that, even without any statistical validation, retweet interactions among the annotated verified users already separate clearly along party and coalition lines, indicating that well-defined political communities exist within this subset.